a homeowner decides to fence in a rectanglar yard for his dog. The dimension of the area to be fenced in are 20ft by 35ft. What is the perimeter of this area. If teh fencing cost $2.19 per foot, how much will it cost him to fence in this area?
1
Expert's answer
2011-10-14T08:09:53-0400
Perimeter=20*2+35*2=110ft
Fence in this area cost: 110*2.19=240.9$